the aptitude test was administered to random sample of 120 SHS students. the sample mean was 104.2 with standard deviation of 12. find 99% confidence interval


1
Expert's answer
2022-05-01T17:29:37-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{s}{\sqrt{n}},"

where:

  • "\bar{x}=104.2" : sample mean
  • zthe chosen z-value, for a 99% confidence interval z = 2.576
  • "s=12" sample standard deviation
  • n = 120 sample size.

So,

"CI=104.2 \pm 2.576\cdot\cfrac{12}{\sqrt{120}}=\\=
104.2\pm2.82=(101.38, 107.02)."